# Mean field equation and Green’s function on the Tate curve

### 16 Mar 2026 v1

@@ Abstract
In this paper, we study the spectrum of the Laplacian on the Tate curve and construct the associated Green’s function as a finite sum, which can be viewed as the non-Archimedean counterpart of the Green’s function on the flat torus in the Archimedean case. And we establish the existence and uniqueness results of the mean field equation on this space. To address this problem, we begin by presenting the solution structure and proving its uniqueness under specific conditions on the finite quotient. We then establish convergence of the solutions to obtain existence results. Notably, the well-posedness of the solution resembles that in the Archimedean case.
@@ 1 Introduction
The mean field equation
@@
arises from statistical mechanics and resembles the Liouville equation, the latter stems from the prescribed Gaussian curvature problem but formulated without a point source term on the right-hand side. A more general formulation is given by the Kazdan-Warner equation. When $\rho=4\pi$, (1.1) is related to the Chern-Simons model [4]. On special domains such as $\Omega\subseteq\mathbb{R}^{n}$ or compact manifolds, the existence and uniqueness of the solution has been studied in [5, 21, 22, 6, 7, 17].
The existence of the solution on graphs has also been investigated. Grigor’yan, Lin, and Yang [9] provided necessary conditions for the existence of solutions on finite graphs, while Huang-Lin-Yau [13] established the existence of solutions on finite weighted graphs. Additionally, Ge and Jiang [8] further proved the existence of solutions on infinite graphs.
Ling and Wang [20, 19] studied the mean field equation on the flat torus $\mathbb{C}/(\mathbb{Z}+\mathbb{Z}\tau)$ for $\tau\in\mathbb{C}\backslash\mathbb{R}$. When $\rho\neq 8m\pi$ with $m\in\mathbb{Z}$, there always exist a solution as the Leray-Schauder degree is nonzero, independent of the torus shape. When $\rho=8\pi$, they demonstrated that the existence and uniqueness of the solution is equivalent to the number of critical points of the Green’s function
$\Delta G(z,w)=\delta_{w}(z)-\frac{1}{V}$ (1.2)
where $V$ denotes the volume of the torus. Specifically, if the mean field equation admits a unique solution, then the Green’s function possesses five critical points, otherwise it has only three.
Since these Liouville-type equations have numerous applications in quantum field theory and string theory [23, 10, 1], and the Green’s function underlies the computation of Veneziano amplitudes. It is natural to investigate a $p$-adic analogue of this problem. Analyses of $p$-adic pseudo-differential operators can be found in [24].
In the genus zero case, Huang-Stoica-Yau-Zhong [15, 16] defined the $p$-adic version of the flat Laplacian via regularizing Vladimirov derivation as
@@
and constructed the associated Green’s function, where $dz$ is the Haar measure on $\mathbb{Q}_{p}$. Besides, in the upper-half plane, Hassan [11] constructed the higher Green’s function.
For genus one, Huang-Rohrlich-Sun-Whyman [14] defined the $p$-adic version of the flat Laplacian on the Tate curve as
$Du(x)=|x|_{p}\int_{\mathbb{Q}_{p}^{\times}/q^{\mathbb{Z}}}\frac{u(z)-u(x)}{|z-x|_{p}^{2}}dz$
where $|q|_{p}<1$. It is easy to check that the operator $D$ is self-adjoint, negative semi-definite and preserves locally constant functions under the inner product induced by the Haar measure $\frac{dx}{|x|_{p}}$ on $\mathbb{Q}_{p}^{\times}/q^{\mathbb{Z}}$. They proved the Green’s function only depends on $d(x,y)=\frac{|x-y|_{p}}{\text{max}\{|x|_{p},|y|_{p}\}}$ and provided a series expression for the Green’s function by correcting the logarithmic singularity on the diagonal.
@@
$\mathbb{Q}_{p}/(\mathbb{Z}+\mathbb{Z}\tau)$ is not a suitable object because $\mathbb{Z}+\mathbb{Z}\tau$ is not closed in $\mathbb{Q}_{p}$, hence the quotient topology is not Hausdorff. In the Archimedean case, we consider the exponential map: $x\mapsto e^{2\pi ix}$, yielding $\mathbb{C}/(\mathbb{Z}+\mathbb{Z}\tau)\rightarrow\mathbb{C}^{\times}/(e^{2\pi i\tau})^{\mathbb{Z}}$. So the Tate curve $\mathbb{Q}_{p}^{\times}/q^{\mathbb{Z}}$ with $|q|_{p}<1$ can be viewed as the non-Archimedean analogy of the flat torus.
@@
Since multiplication by a unit in $\mathbb{Z}_{p}^{*}$ leaves $D$ invariant, we may restrict to the case $q=p^{m}$, then we choose the fundamental domain as $\cup_{k=0}^{m-1}p^{l}\mathbb{Z}_{p}^{*}$. One can verify that the definition is independent of the choice of the fundamental domain.
@@
$DG(x,y)=\delta_{y}(x)-\frac{1}{V}\ \text{in}\ \mathbb{Q}_{p}^{\times}/q^{\mathbb{Z}}$ (1.3)
$Du+\rho e^{u}=\rho\delta_{y}\ \text{in}\ \mathbb{Q}_{p}^{\times}/q^{\mathbb{Z}}$ (1.4)
A standard approach to constructing the Green’s function employs the spectrum of the operator $D$ and its eigenfunctions. When $m=1$, the fundamental domain is a locally compact group, allowing us to obtain its spectrum and eigenfunctions via characters, some of these works have been established in [12]. We extend the results to the general $m$ case through zero extension although the operator $D$ is a nonlocal operator, and make minor modifications in the case when the eigenvalue is zero. Using orthogonality of the characters, we express the Green’s function as a finite sum, the resulting formula differs form that in [14].
Let $\boldsymbol{L}=\boldsymbol{P}-\text{diag}(\boldsymbol{P}\boldsymbol{1}_{m})$ and $\boldsymbol{R}=-(1-p^{-1})\boldsymbol{L}$, where $(\boldsymbol{P})_{ij}=p^{-|i-j|}$, so we have orthogonal matrix $\boldsymbol{Q}$ to diagonalize $\boldsymbol{R}$. And we define $\phi_{h,c,n}(x)=e^{\frac{2\pi ih}{p-1}\text{ord}(a_{0})}e^{2\pi i\{\frac{c\text{ln}\frac{x}{a_{0}}}{p^{n}}\}_{p}}$, where $x=a_{0}+a_{1}p+\dots\in\mathbb{Z}_{p}^{*}$, $0\leq h\leq p-2$ and $c\in\mathbb{Z}_{p}^{*}$. Then our first main result is:
@@
Inspired by the existence results for the Green’s function [14], we first consider the mean field equation on the finite quotient $\mathbb{Q}_{p}^{\times}/p^{m\mathbb{Z}}/(1+p^{d}\mathbb{Z}_{p})$. Since there are finitely many points in the finite quotient, we can solve the equation by numerical experiments. Based on observations from numerical results, we investigate a universal upper bound of the solution and establish uniqueness when the parameter $\rho\leq C(p,m)$. We then study the convergence of the solutions to obtain existence on $\mathbb{Q}_{p}^{\times}/p^{m\mathbb{Z}}$. Similar to the arguments on the finite quotient, the uniqueness can be preserved, yielding a unique solution. The optimal upper bound of $\rho$ can be viewed as analogous to the $8\pi$ threshold in Archimedean case. Then our second main result is:
@@
When $\rho\leq\begin{cases}(1-2p^{-1})(1-p^{-1})(\delta_{m,1}(1+\lambda_{1})-\lambda_{1}),\ &p>2\\ \frac{1}{8}(\delta_{m,1}(1+\lambda_{1})-\lambda_{1}),&p=2\end{cases}$, then there exists a unique solution of the mean field equation (1.4). Where $\lambda_{1}$ is the smallest nonzero eigenvalue of $\boldsymbol{L}$.
Our ultimate goal is to connect the Green’s function with the mean field equation to explore the critical point properties of the Green’s function. Since the normal derivative on the Green’s function is trivial due to its local constancy, our first step will be to define a suitable derivative on the Green’s function in future work.
@@ 1.1 Outline of the paper
In section 3, we study the mean field equation on finite quotients of the Tate curve. We derive a uniform positive upper bound $C_{p}$(depending only on $p$) for the solution and demonstrate its central symmetry, cyclic symmetry, and monotonicity. Furthermore, we prove that the solution is radial symmetric and dependents only on $|x|_{p}$, $|y|_{p}$ and $|x-y|_{p}$ when $\rho\leq C_{p}(1-p^{-m})$, which is similar to the property of the Green’s function. Lastly we establish uniqueness for $\rho\leq C_{p}\mu_{1}$, where $\mu_{1}$ is the smallest nonzero eigenvalue of D-D.
@@ 2 Spectrum of D and the Green’s function
This section focuses on the spectrum of $D$ and its eigenfunctions. First, we consider the case $m=1$, then the fundamental domain is $\mathbb{Z}_{p}^{*}$, which is a compact multiplicative group. We analyze $D$ via characters of $\mathbb{Z}_{p}^{*}$. After that, we construct the Green’s function with a finite sum. Lastly, we generalize it to the general $m$ case.
@@ 2.1 Multiplicative Characters of $\mathbb{Z}_{p}^{*}$
Let $g$ be a primitive root of $(\mathbb{Z}/p\mathbb{Z})^{\times}$, as the dual group of $(\mathbb{Z}/p\mathbb{Z})^{\times}$ is isomorphic to itself, so for $m,h\in\{0,1,\dots,p-2\}$, the characters for $(\mathbb{Z}/p\mathbb{Z})^{\times}$ are given by
$\chi_{h}(g^{m})=e^{\frac{2\pi ih}{p-1}m}$
As the dual group of $\mathbb{Z}_{p}$ is isomorphic to $\mathbb{Q}_{p}/\mathbb{Z}_{p}$, so the characters on $\mathbb{Z}_{p}$ are $\chi_{t}(x)=e^{2\pi i\{tx\}_{p}}$, where $x\in\mathbb{Z}_{p},t\in\mathbb{Q}_{p}/\mathbb{Z}_{p}$. And we notice that there is a continuous isomorphism $\mathbb{Z}_{p}\cong 1+p\mathbb{Z}_{p},x\mapsto e^{px}$. So the characters of $1+p\mathbb{Z}_{p}$ can be pulled back from the isomorphism as
$\chi_{t}(x)=e^{2\pi i\{t\frac{\text{ln}x}{p}\}_{p}}$
where $x\in 1+p\mathbb{Z}_{p},t\in\mathbb{Q}_{p}/\mathbb{Z}_{p}$. Denoting $t=a_{-n}p^{-n}+a_{-n+1}p^{-n+1}+\dots+a_{-1}p^{-1}=\frac{a_{-n}+a_{-n+1}p+\dots+a_{-1}p^{n-1}}{p^{n}}$, and taking $c=a_{-n}+a_{-n+1}p+\dots+a_{-1}p^{n-1}$, we rewrite the character $\chi_{t}(x)$ as
$\chi_{c,n+1}(x)=e^{2\pi i\{\frac{c\text{ln}x}{p^{n+1}}\}_{p}}$
So $\forall x=a_{0}+a_{1}p+\dots\in\mathbb{Z}_{p}^{*}$, the character of $\mathbb{Z}_{p}^{*}$ is trivial character $1$ and
$\chi_{h,c,n}(x)=\chi_{h}(a_{0})\chi_{c,n}(\frac{x}{a_{0}})=e^{\frac{2\pi ih}{p-1}\text{ord}(a_{0})}e^{2\pi i\{\frac{c\text{ln}\frac{x}{a_{0}}}{p^{n}}\}_{p}}$ (2.1)
where $h\in\{0,1,\dots,p-2\},n\in\mathbb{N_{+}}$. And the conductor of $\chi_{h,c,n}$ is $n$.
@@ 2.2 Eigenvalues and eigenfunctions for $D$ when $m=1$
$=\int_{\mathbb{Z}_{p}^{*}}e^{\frac{2\pi i(h-h^{\prime})}{p-1}\text{ord}(a_{0})}e^{2\pi i\{(\frac{c}{p^{n}}-\frac{c^{\prime}}{p^{n^{\prime}}})\text{ln}\frac{x}{a_{0}}\}_{p}}dx$
$=\sum_{a_{0}=1}^{p-1}e^{\frac{2\pi i(h-h^{\prime})}{p-1}\text{ord}(a_{0})}\int_{1+p\mathbb{Z}_{p}}e^{2\pi i\{(\frac{c}{p^{n}}-\frac{c^{\prime}}{p^{n^{\prime}}})\text{ln}\frac{x}{a_{0}}\}_{p}}dx$
$=p^{-1}\sum_{a_{0}=1}^{p-1}e^{\frac{2\pi i(h-h^{\prime})}{p-1}\text{ord}(a_{0})}\int_{\mathbb{Z}_{p}}e^{2\pi i\{(\frac{c}{p^{n}}-\frac{c^{\prime}}{p^{n^{\prime}}})y\}_{p}}dy$
$=p^{-1}\sum_{a_{0}=1}^{p-1}e^{\frac{2\pi i(h-h^{\prime})}{p-1}\text{ord}(a_{0})}\boldsymbol{1}_{\mathbb{Z}_{p}}(\frac{c}{p^{n}}-\frac{c^{\prime}}{p^{n^{\prime}}})$
$=(1-p^{-1})\delta_{h,c,n}^{h^{\prime},c^{\prime},n^{\prime}}$
@@
Since the conductor of $\chi_{h,c,n}$ is $n$, it is trivial on $1+p^{n}\mathbb{Z}_{p}$ but nontrivial on $1+p^{n-1}\mathbb{Z}_{p}$. By the first orthogonality relation of the character, $\forall s\in\{1,2,\dots,n-1\}$, we have
$\sum_{x\in(1+p^{s}\mathbb{Z}_{p})/(1+p^{n}\mathbb{Z}_{p})}\chi_{h,c,n}(x)=0$
@@
$f(x)=\sum_{\chi_{\xi}\in\text{Char}(\mathbb{Z}_{p}^{*})}\hat{f}(\xi)\chi_{\xi}(x)$
@@
$=\lambda_{f}\sum_{\chi_{\xi}\in\text{Char}(\mathbb{Z}_{p}^{*})}\hat{f}(\xi)\chi_{\xi}(x)$
@@
$\sum_{\chi_{\xi}\in\text{Char}(\mathbb{Z}_{p}^{*})}(\lambda_{\xi}-\lambda_{f})\hat{f}(\xi)\chi_{\xi}(x)=0$

### 08 Apr 2026 v2

# The Mean field equation on the Tate curve

@@ Abstract
In this paper, we study the spectrum of the Laplacian on the Tate curve and construct the associated Green’s function as a finite sum, which can be viewed as the non-Archimedean counterpart of the Green’s function on the flat torus in the Archimedean case. Moreover, we establish existence and uniqueness results of the mean field equation on this space. To address the problem, we first prove the structure of solutions on finite quotients, and prove the existence on the Tate curve by the convergence of such solutions. We also prove the uniqueness of the solutions for some parameter region. Notably, the well-posedness of the solution resembles that in the Archimedean case.
@@ 1 Introduction
The mean field equation,
@@
rooted in statistical mechanics, is formally analogous to the Liouville equation; however, the latter originates from the prescribed Gaussian curvature problem and is typically formulated without a point source term on the right-hand side. A more general formulation is given by the Kazdan-Warner equation [17]. When $\rho=4\pi$, (1.1) is related to the Chern-Simons model [4]. On special domains such as $\Omega\subseteq\mathbb{R}^{n}$ or compact manifolds, the existence and uniqueness of the solution have been studied in [5, 23, 24, 6, 7, 18].
The existence of solutions to these type of equations on graphs has been investigated. Grigor’yan, Lin, and Yang [9] provided various conditions for the existence of the Kazdan-Warner equation on finite graphs, Ge and Jiang [8] further proved the existence of solutions on infinite graphs.
Huang-Lin-Yau [13] established the existence of the mean field equation on finite weighted graphs via the calculus of variations. Further Lin and Yang introduced a mean-field type heat flow on finite graphs [25]. In the same year, Liu studied the mean-field equation with a positive prescribed function by using the method of Brouwer degree [27]. Later, Sun and Wang promoted their results to the sign-changing prescribed function [28]. After that, Liu and Zhang further developed the heat flow method and resolved the convergence problem for the mean-field equation with a sign-changing prescribed function [26]. In recent years, Li and Zhang systematically investigated the mean-field equation with a linear term by means of variational methods [19]. And Liu extended the results to nonlocal operators via considering the fractional Laplacian $(-\Delta)^{s}$, this work unifies and generalizes the results of [25, 28, 26] for the case $s=1$.
Lin and Wang [22, 21] studied the mean field equation on the flat torus $\mathbb{C}/(\mathbb{Z}+\mathbb{Z}\tau)$ for $\tau\in\mathbb{C}\backslash\mathbb{R}$. When $\rho\neq 8m\pi$ for $m\in\mathbb{Z}$, there always exists a solution as the Leray-Schauder degree is nonzero, which is independent of the torus shape $\tau$. When $\rho=8\pi$, they demonstrated that the existence and uniqueness of the solution depend on the number of critical points of the Green’s function
$\Delta G(z,w)=\delta_{w}(z)-\frac{1}{V},$ (1.2)
where $V$ denotes the volume of the torus. In particular, the mean field equation admits a unique solution if and only if the Green’s function possesses five critical points (otherwise it has only three).
Since these Liouville-type equations have numerous applications in quantum field theory and string theory [29, 10, 1], and the Green’s function underlies the computation of Veneziano amplitudes. It is natural to investigate a $p$-adic analogue of this problem. Analysis of $p$-adic pseudo-differential operators can be found in [30].
In the genus-zero case, Huang-Stoica-Yau-Zhong [15, 16] defined the $p$-adic version of the flat Laplacian via regularizing Vladimirov derivation as
@@
and constructed the associated Green’s function, where $dz$ is the Haar measure on the $p$-adic field $\mathbb{Q}_{p}$. Besides, in the upper-half plane, Hassan [11] constructed the higher Green’s function.
For the genus-one case, Huang-Rohrlich-Sun-Whyman [14] defined the $p$-adic version of the flat Laplacian on the Tate curve as
$Du(x)=|x|_{p}\int_{\mathbb{Q}_{p}^{\times}/q^{\mathbb{Z}}}\frac{u(z)-u(x)}{|z-x|_{p}^{2}}dz,$
where $|q|_{p}<1$. One easily checks that the operator $D$ is self-adjoint and negative semi-definite under the inner product induced by the Haar measure $\frac{dx}{|x|_{p}}$ on $\mathbb{Q}_{p}^{\times}/q^{\mathbb{Z}}$, which preserves locally constant functions. They proved the Green’s function only depends on $d(x,y)=\frac{|x-y|_{p}}{\max\{|x|_{p},|y|_{p}\}}$ and provided a series expression for the Green’s function by correcting the logarithmic singularity on the diagonal.
@@
$\mathbb{Q}_{p}/(\mathbb{Z}+\mathbb{Z}\tau)$ is not a suitable object because $\mathbb{Z}+\mathbb{Z}\tau$ is not closed in $\mathbb{Q}_{p}$, hence the quotient topology is not Hausdorff. In the Archimedean case, we consider the exponential map: $x\mapsto e^{2\pi ix}$, yielding $\mathbb{C}/(\mathbb{Z}+\mathbb{Z}\tau)\rightarrow\mathbb{C}^{\times}/(e^{2\pi i\tau})^{\mathbb{Z}}$. Based on that, the Tate curve $\mathbb{Q}_{p}^{\times}/q^{\mathbb{Z}}$ with $|q|_{p}<1$ can be viewed as the non-Archimedean analogy of the flat torus.
@@
Since multiplication by a unit in $\mathbb{Z}_{p}^{*}$ leaves $D$ invariant, we may restrict to the case $q=p^{m}$, and choose the fundamental domain as $\bigcup_{k=0}^{m-1}p^{k}\mathbb{Z}_{p}^{*}$. One can verify that the definition is independent of the choice of the fundamental domain.
@@
$DG(x,y)=\delta_{y}(x)-\frac{1}{V}\ \text{in}\ \mathbb{Q}_{p}^{\times}/q^{\mathbb{Z}},$ (1.3)
$Du+\rho e^{u}=\rho\delta_{y}\ \text{in}\ \mathbb{Q}_{p}^{\times}/q^{\mathbb{Z}}.$ (1.4)
A standard approach to construct the Green’s function employs the spectrum of the operator $D$ and its eigenfunctions. When $m=1$, the fundamental domain is a locally compact group, allowing us to obtain its spectrum and eigenfunctions via characters, some of these works have been established in [12]. We extend the results to the general $m$ case through zero extension although the operator $D$ is a nonlocal operator, and make minor modifications in the case when the eigenvalue is zero. Using orthogonality of the characters, we express the Green’s function as a finite sum, the resulting formula differs form that in [14].
Letting $\boldsymbol{R}=(1-p^{-1})(\boldsymbol{P}-\text{diag}(\boldsymbol{P}\boldsymbol{1}_{m}))$, where $(\boldsymbol{P})_{ij}=p^{-|i-j|}$ and $\text{diag}(\boldsymbol{v})$ denotes the diagonal matrix with the vector $\boldsymbol{v}$ as its diagonal entries. We have orthogonal matrix $\boldsymbol{Q}$ to diagonalize $\boldsymbol{R}$, and we define
$\phi_{h,c,n}(x)=e^{\frac{2\pi ih}{p-1}\text{ord}(a_{0})}e^{2\pi i\left\{\frac{c\ln\left(\frac{x}{a_{0}}\right)}{p^{n}}\right\}_{p}},$
where $x=a_{0}+a_{1}p+\dots\in\mathbb{Z}_{p}^{*}$, $0\leq h\leq p-2$ and $c\in\mathbb{Z}_{p}^{*}$ with $cp^{n-1}\in\mathbb{Q}_{p}/\mathbb{Z}_{p}$. Then our first main result is as follows:
@@
Inspired by the existence results for the Green’s function [14], we consider the mean field equation on the finite quotient $\mathbb{Q}_{p}^{\times}/p^{m\mathbb{Z}}/(1+p^{d}\mathbb{Z}_{p})$ first. The operator $D$ on this finite quotient simplifies to a graph Laplacian, reducing the problem to an equation on a finite weighted graph. Motivated by these, we prove a universal upper bound for the solution and establish its uniqueness under the condition of the parameter $\rho$. We then study the convergence of the solutions to obtain the existence result on $\mathbb{Q}_{p}^{\times}/p^{m\mathbb{Z}}$. Similar to the arguments on the finite quotient, the uniqueness can be also derived. The optimal upper bound of $\rho$ can be viewed as analogous to the $8\pi$ threshold in Archimedean case. Then our second main result is as follows:
@@
When $\rho\in(0,\rho^{*})$, then there exists a unique solution of the mean field equation (1.4), where $\rho^{*}=\theta_{p,m}e^{-M}$ with
[…]
and $\theta_{p,m}$ is the smallest nonzero eigenvalue of D-D. Furthermore, the solution is radially symmetric and strictly monotone.
Our goal is to connect the Green’s function with the mean field equation to explore the critical point properties of the Green’s function in the future works. Since the results in Archimedean case were obtained indirectly through the mean-field equation, our results will provide some preliminary steps toward this problem.
@@ 1.1 Outline of the paper
In section 3, we study the mean field equation on finite quotients of the Tate curve. We derive a uniform positive upper bound $M$ (depending only on $p$ and $m$) for the solution and demonstrate its central symmetry, cyclic symmetry, and monotonicity. Furthermore, we prove that the solution is radially symmetric and dependents only on $|x|_{p}$, $|y|_{p}$ and $|x-y|_{p}$ when $\rho\leq(1-p^{-m})e^{-M}$, which is similar to the property of the Green’s function. Lastly we establish uniqueness for $\rho<\rho^{*}$.
@@ 2 Spectrum of D and the Green’s function
This section focuses on the spectrum of $D$ and its eigenfunctions. First, we consider the case $m=1$, where the fundamental domain is $\mathbb{Z}_{p}^{*}$, which is a compact multiplicative group. We analyse $D$ via characters of $\mathbb{Z}_{p}^{*}$ and construct the Green’s function with a finite sum. Then we generalize it to the general $m$ case through zero extension.
@@ 2.1 Multiplicative Characters of $\mathbb{Z}_{p}^{*}$
Let $g$ be a primitive root of $(\mathbb{Z}/p\mathbb{Z})^{\times}$, as the dual group of $(\mathbb{Z}/p\mathbb{Z})^{\times}$ is isomorphic to itself, so for $w,h\in\{0,1,\dots,p-2\}$, the characters for $(\mathbb{Z}/p\mathbb{Z})^{\times}$ are given by
$\chi_{h}(g^{w})=e^{\frac{2\pi ih}{p-1}w}.$
As the dual group of $\mathbb{Z}_{p}$ is isomorphic to $\mathbb{Q}_{p}/\mathbb{Z}_{p}$, so the characters on $\mathbb{Z}_{p}$ are $\chi_{t}(x)=e^{2\pi i\{tx\}_{p}}$, where $x\in\mathbb{Z}_{p},t\in\mathbb{Q}_{p}/\mathbb{Z}_{p}$. And we notice that there is a continuous isomorphism
$\mathbb{Z}_{p}$ $\cong 1+p\mathbb{Z}_{p}$
$x$ $\mapsto e^{px}.$
So the characters of $1+p\mathbb{Z}_{p}$ can be pulled back from the isomorphism as
$\chi_{t}(x)=e^{2\pi i\left\{t\frac{\ln x}{p}\right\}_{p}},$
where $x\in 1+p\mathbb{Z}_{p},t\in\mathbb{Q}_{p}/\mathbb{Z}_{p}$ and $\{\cdot\}_{p}$ denotes the $p$-adic fractional part. Letting
$t=c_{-n}p^{-n}+c_{-n+1}p^{-n+1}+\dots+c_{-1}p^{-1}=\frac{c_{-n}+c_{-n+1}p+\dots+c_{-1}p^{n-1}}{p^{n}}$
and taking
$c=c_{-n}+c_{-n+1}p+\dots+c_{-1}p^{n-1},$
we rewrite the character $\chi_{t}(x)$ as
$\chi_{c,n+1}(x)=e^{2\pi i\left\{\frac{c\ln x}{p^{n+1}}\right\}_{p}}.$
So $\forall x=a_{0}+a_{1}p+\dots\in\mathbb{Z}_{p}^{*}$ for $a_{0}\neq 0$, the characters of $\mathbb{Z}_{p}^{*}$ are trivial character $1$ and
$\chi_{h,c,n}(x)=\chi_{h}(a_{0})\chi_{c,n}\left(\frac{x}{a_{0}}\right)=e^{\frac{2\pi ih}{p-1}\text{ord}(a_{0})}e^{2\pi i\left\{\frac{c\ln\left(\frac{x}{a_{0}}\right)}{p^{n}}\right\}_{p}},$ (2.1)
where $h\in\{0,1,\dots,p-2\},n\in\mathbb{N_{+}}$. From (2.1) one can easily check that the conductor of the character $\chi_{h,c,n}$ is $n$.
@@ 2.2 Eigenvalues and eigenfunctions for $D$ when $m=1$
$=\int_{\mathbb{Z}_{p}^{*}}e^{\frac{2\pi i(h-h^{\prime})}{p-1}\text{ord}(a_{0})}e^{2\pi i\left\{\left(\frac{c}{p^{n}}-\frac{c^{\prime}}{p^{n^{\prime}}}\right)\ln\left(\frac{x}{a_{0}}\right)\right\}_{p}}dx$
$=\sum_{a_{0}=1}^{p-1}e^{\frac{2\pi i(h-h^{\prime})}{p-1}\text{ord}(a_{0})}\int_{1+p\mathbb{Z}_{p}}e^{2\pi i\left\{\left(\frac{c}{p^{n}}-\frac{c^{\prime}}{p^{n^{\prime}}}\right)\ln\left(\frac{x}{a_{0}}\right)\right\}_{p}}dx$
$=p^{-1}\sum_{a_{0}=1}^{p-1}e^{\frac{2\pi i(h-h^{\prime})}{p-1}\text{ord}(a_{0})}\int_{\mathbb{Z}_{p}}e^{2\pi i\left\{\left(\frac{c}{p^{n}}-\frac{c^{\prime}}{p^{n^{\prime}}}\right)y\right\}_{p}}dy$
$=p^{-1}\sum_{a_{0}=1}^{p-1}e^{\frac{2\pi i(h-h^{\prime})}{p-1}\text{ord}(a_{0})}\boldsymbol{1}_{\mathbb{Z}_{p}}\left(\frac{c}{p^{n}}-\frac{c^{\prime}}{p^{n^{\prime}}}\right)$
$=(1-p^{-1})\delta_{h,c,n}^{h^{\prime},c^{\prime},n^{\prime}},$
@@
[…]
$\sum_{x\in(1+p^{s}\mathbb{Z}_{p})/(1+p^{n}\mathbb{Z}_{p})}\chi_{h,c,n}(x)=0,$
@@
$f(x)=\sum_{\chi_{\xi}\in\text{Char}(\mathbb{Z}_{p}^{*})}\hat{f}(\xi)\chi_{\xi}(x),$
@@
$=\lambda_{f}\sum_{\chi_{\xi}\in\text{Char}(\mathbb{Z}_{p}^{*})}\hat{f}(\xi)\chi_{\xi}(x),$
@@
$\sum_{\chi_{\xi}\in\text{Char}(\mathbb{Z}_{p}^{*})}(\lambda_{\xi}-\lambda_{f})\hat{f}(\xi)\chi_{\xi}(x)=0.$